OpenStudy (anonymous):

Use the substitution method to find the solutions to the system of equations. y=-x y=x^2-3

OpenStudy (saifoo.khan):

-x = x^2 - 3 We can equal Y's.

OpenStudy (rogue):

y=-x y=x^2-3 y = y -x = x^2 - 3 0 = x^2 + x - 3 Use the quadratic formula.\[x = \frac {-b \pm \sqrt {b^2 - 4ac}}{2a}\]\[x = \frac {-1 \pm \sqrt {1 + 12}}{2}\]\[x = \frac {1}{2}(-1 - \sqrt {13}), x = \frac {1}{2} (-1 + \sqrt {13})\]

OpenStudy (anonymous):

thank you